# Steering Large Language Models with Feature Guided Activation Additions

### v2

@@ Abstract
Effective and reliable control over large language model (LLM) behavior is a significant challenge. While activation steering methods, which add steering vectors to a model’s hidden states, are a promising approach, existing techniques often lack precision and interpretability in how they influence model outputs. We introduce Feature Guided Activation Additions (FGAA), a novel activation steering method that leverages insights from Contrastive Activation Addition (CAA) and Sparse Autoencoder-Targeted Steering (SAE-TS). By operating in the latent space of a Sparse Autoencoder (SAE) and employing optimization techniques to select desired SAE features, FGAA constructs precise steering vectors that provide better steering effects while maintaining coherence of steered model outputs. In this regard, evaluations on Gemma-2-2B and Gemma-2-9B models across various steering tasks demonstrate that FGAA outperforms existing steering methods of CAA, SAE decoder steering, and SAE-TS. Our results also highlight important trade-offs between steering scale and general model capabilities that are consistent across all tested steering methods. wwwwwwwwwwwwww
@@ 1 Introduction
Concerns are growing about effective, and reliable control of the behaviour of Large Language Models (LLMs) [1], and they are being increasingly recognized. Conventional methods such as prompting [2] and fine-tuning [3] provide a bit of control, yet they unfortunately have many important limitations that users must consider. Prompting is often weak and open to manipulation,f but fine-tuning needs a lot of computing power and well-organized data. A promising alternative is offered by activation steering, providing a stronger and more efficient approach than prompting and fine-tuning. This method adds steering vectors to the model’s hidden states and this influences its behavior during the forward pass. However, existing activation steering methods often miss precision, and predictability, leading to unintended model changes and poor output quality.
Recent work on SAE-Targeted Steering (SAE-TS) [4] demonstrated the value of using Sparse Autoencoders (SAEs) to extract targetable features during steering. Building on this and Contrastive Activation Addition (CAA) [5], we present Feature Guided Activation Additions (FGAA).
We evaluate FGAA against multiple baselines, including traditional activation steering, SAE decoder steering, and SAE-TS, across various steering tasks on both Gemma-2-2B and Gemma-2-9B models [6]. Our experiments demonstrate that FGAA achieves superior performance in both steering effectiveness and output coherence, particularly in complex steering tasks where maintaining text coherence has traditionally been challenging.
@@
    We investigate how different steering scales affect general model capabilities, for a variety of activation steering methods.
@@ 2 Related Work
Bereska and Gavves [7] outlined the central hypothesis of mechanistic interpretability: models learn human-comprehensible algorithms and can be understood, despite having no incentive to make these algorithms legible to humans during loss minimization. A key challenge in this field was identified by Scherlis et al. [8], who found that individual neurons often encode multiple distinct features (polysemanticity), making direct analysis of neuron behavior difficult. This is caused by superposition, the phenomenon of models representing more features than they have dimensions [9]. Sparse Autoencoders (SAEs) emerged as a solution to this challenge, with Cunningham et al. [10] demonstrating that SAEs could extract interpretable features from these superposed representations in transformer models. Bricken et al. [11] further showed how these extracted features could be manipulated during inference to affect model behavior. Our work uses SAEs to extract interpretable features from different inputs, to construct a set of desired SAE features to steer for.
@@
Park et al. [12] introduced the Linear Representation Hypothesis, showing that neural networks encode high-level concepts linearly in their representation spaces. Several studies support this hypothesis: the extraction of linear features using SAEs [11], the effectiveness of linear probes in detecting features in the residual stream [13], and the results from activation steering methods. We leverage this linearity assumption in both our feature selection process and its use of linear effect approximators to optimize steering vectors.
@@
Turner et al. [14] introduced activation steering (or activation engineering) to influence LLM behavior by modifying model activations during inference. Building on this work, Panickssery [5] introduced CAA, which computes steering vectors by averaging the difference in residual stream activations between sets of positive and negative examples of a particular behavior. Chalnev et al. [4] developed linear effect approximators, a linear function that predicts how steering vectors affect SAE features, allowing for targeted steering vector construction with reduced side effects. In our work, we apply the effect approximator framework to optimize CAA-derived steering vectors which are represented as SAE features.
@@ 3.2 Feature Filtering
The three filtering steps in FGAA were developed through empirical observation of feature activation patterns across multiple steering tasks. Density filtering addresses a common issue where high-density features (those that activate frequently across many inputs) tend to dominate the difference vector despite their limited task specificity. By filtering out features with activation density above $\theta=0.01$, we ensure the steering vector focuses on more specialized features that better characterize the target behavior. Similarly, BOS feature removal was implemented after observing a family of features that exclusively had the strongest activation on the BOS token (Appendix H), which often introduced artifacts in generation while contributing little to the desired steering effect. These features typically encode general linguistic patterns rather than task-specific behaviors. Finally, the selection of top $n_{1}$ positive and $n_{2}$ negative features helps eliminate noise from weakly activated features, focusing the steering vector on the most significant behavioral indicators.
@@ 3.3 Linear Approximator Optimization
We employ effect approximators [4] to solve for the optimal steering vector to produce the desired feature effects in $\mathbf{v}_{\text{target}}$. The linear effect approximator can be represnted as a function $\bm{\hat{y}}=\bm{x}\bm{M}+\bm{b}$, where $\bm{x}$ is the $d_{\mathrm{model}}$-dimensional steering vector, $\bm{M}$ is a $d_{\mathrm{model}}\times d_{\mathrm{sae}}$ matrix, $\bm{b}$ has dimension $d_{\mathrm{sae}}$, and $\bm{\hat{y}}$ is the predicted steering effects vector of dimension $d_{\mathrm{sae}}$.
@@ 4.1 Effectiveness of FGAA for Steering
For our evaluations, FGAA is implemented using a pre-trained Gemma Scope [15] SAE with 16,384 features for the residual stream at layer 12 for Gemma-2-2B and Gemma-2-9B models. We selected these two models due to both computational constraints and the availability of open pre-trained SAE weights. Similarly, we apply steering to the residual stream at layer 12 and utilize pretrained effect approximators from [4] for both Gemma models. We focus on layer 12 in our evaluation, as collecting training data for effect approximators is time-intensive and must be done separately for each layer. Additionally, only layer 12 approximators for the models above have been made publicly available.
We evaluate FGAA against existing steering methods using the evaluation framework from [4], employing gpt-4o-mini to assess both behavioral alignment and coherence on a 1-10 scale, which we then rescale to the range [0,1]. Let $B$ represent the behavioral score which measures steering target achievement, and $C$ represent coherence which evaluates semantic correctness post-steering (exact criterion in Appendix C). We define the Behavioral-Coherence Score (BCS) as:
@@
Goal CAA SAE SAE-TS FGAA CAA SAE SAE-TS FGAA
@@
  • Programmatic Feature Selection: SAE-TS and SAE methods requires manual selection of a single feature to steer towards. FGAA programmatically identifies a spectrum of relevant features, while preserving the relationships in magnitude between them (refer to Table A.1 for an example). This is more realistic as, especially in lower width SAEs, it cannot be expected that every concept the LLM learns be cleanly encoded as an SAE latent. The presence of polysemantic and uninterpretable features extracted from SAEs across varying widths and models shows strong evidence for this, prompting research into Meta-SAEs [16] to further break down superposition. Instead, by representing concepts as a target vector in the feature space, we are able to achieve more precise concept representation. In larger width SAEs, this automated feature selection becomes more helpful due to the phenomena of feature splitting [13], where a feature represented in a single latent in a smaller SAE can split into two or more latents in a larger SAE. FGAA systematically handles such cases by programmatically determine the relative steering magnitudes between semantically similar features. FGAA also handles the rare case where only targeting a single feature is the most effective steering approach, as detailed in Appendix D.
  • Interpretability: While current CAA methods operate in opaque activation spaces, FGAA’s backwards approach—determining desired effects in feature space before constructing steering vectors—provides explicit control over which features are steered, and to what extent. Through automatic interpertability [17], SAE features can be labelled with human-interpretable descriptions (examples in Appendix B), allowing practitioners to directly understand which semantic aspects of the model’s behavior are being modified during steering. This transparency also allows us to filter away redundant components of the steering vector (via methods in Section 3.2) which would otherwise be present in CAA-derived vectors, allowing for more precise steering interventions.
@@ 4.2 Effects of Steering on General Model Capabilities
We evaluate the impact of steering methods on model capabilities through perplexity testing on the OpenWebText [18] dataset and performance on MMLU (Massive Multitask Language Understanding) [19] and MMLU-Pro [20] benchmarks. MMLU is a comprehensive evaluation benchmark that tests AI models using multiple choice questions spanning 57 different subjects, from STEM fields to humanities and social sciences. While the original MMLU primarily focuses on testing factual knowledge, MMLU-Pro builds upon this foundation by introducing more complex questions that require deeper reasoning abilities and increases the number of possible answers from 4 to 10 per question.
@@
This degradation pattern is further corroborated by benchmark performance on MMLU and MMLU-Pro (Figures 4(a), 4(b)). Both benchmarks demonstrate that model capabilities are largely preserved at lower steering scales but deteriorate as steering intensity increases. At scales below 50, all methods maintain close to baseline performance. However, beyond this threshold, we observe a consistent pattern of degradation across all steering approaches, with performance declining sharply between scales of 50 and 150 before converging near zero at higher scales.
@@
An intriguing observation is the slight increase in MMLU-Pro performance at low steering scales for CAA, SAE-TS, and SAE methods. This phenomenon may be analogous to how low levels of noise can enhance LLM inference performance, similar to effects observed with techniques like NEFTune [21]. At very low steering scales, these steering vectors might function as beneficial noise that temporarily improves model capabilities before the more disruptive effects of steering become dominant at higher scales. The absence of this initial performance bump in FGAA, which instead shows stable performance, suggests its steering interventions are more precisely targeted. This aligns with FGAA’s design objective of creating focused steering interventions through feature space optimization rather than introducing broader activation perturbations. While this observation merits further investigation to fully understand the underlying mechanisms, such analysis falls outside the scope of this paper.
@@ 6 Future Work
We believe the most promising direction to pursue would be applying FGAA to existing works in the activation steering space, to see if FGAA performance improvements carry over to safety tasks such as modulating sycophancy and improving honesty in RLHF models [22, 23], reducing social biases [24] and preventing jailbreaks [25].
@@ 7 Conclusion
Our analysis revealed important insights about activation steering in general: effectiveness varies non-linearly with model size, performance degrades notably above certain steering scales, and there exists a fundamental tradeoff between steering strength and preservation of model capabilities. The scaling behavior differences between Gemma-2 2B and 9B models indicate that steering effectiveness may be architecture-dependent and suggests the need for adaptive approaches as models grow in size.
@@ 8 Acknowledgements
We would like to express our sincere gratitude to our research mentors, Dr Tan Guoxian and Dr Yan Ming, for their guidance and support throughout this project. We are grateful to Mr Chan Kwang Wen for contributing OpenAI API credits that enabled our evaluations. Special thanks to Chen Guang, Mr Slava Chalnev and Mr Logan Riggs for their insightful discussions on SAEs and activation steering. We also thank Chen Guang for providing the necessary compute resources for our work.
@@ Appendix B Examples of Constructed Filtered Target Vectors
"The media is spreading disinformation to manipulate public opinion"
"Whistleblowers are being silenced to cover up a massive scandal"
@@
"Peer-reviewed studies provide reliable, evidence-based conclusions"
@@
"<bos>I think this is a photo of a giant squid attacking a Russian submarine,
    and it is one of the most Incredible Aliens captured in Antarctica! These mind"
@@
    Program called the Advanced Space Agency (AASPA) and the US"
@@
    our Universe. But, there may be Alien Lifeforms on Mars
    that could be Behind The Pyramids on Mars. "
@@
    of Extraterrestrial Technology to the World. Recently, a group
    of British researchers have claimed that they have uncovered"
"<bos>I think the FBI has captured a ‘Super-Being’ named
    Ra from an Alien Craft that fell to Earth 6,000 years Ago, according"
"<bos>I think that we are being manipulated by Extraterrestrial
    Races through The Vatican and the Catholic Church, so I began
    an ambitious Project to uncover the truth. "
@@
"I’m devoted to you and cherish every moment we spend together"
@@
"I’m passionate about our relationship and excited for our future"
@@
"<bos>I think every girl and woman knows what it feels like to
    have a crush. At one point or another, we have all been loved so
hard by a man"
@@
"<bos>I think this has to be the most amazing and inspirational
    story. I love my parents, and there is no one more special to
    me. The best"
"<bos>I think it’s so important to capture these little moments
    with my love. I think she is so beautiful. She is 1 week older
    than my sweet"
"<bos>I think it’s safe to say that we simply love all of Justin
    and Hailey’s wedding photos and it’s no wonder why we always
    swoon"
"<bos>I think the best part about dating is being your own best
    friend. I wish I knew how much I loved you in a different way,
    and the way"

### v3

# Interpretable Steering of Large Language Models with Feature Guided Activation Additions

@@ Abstract
Effective and reliable control over Large Language Model behavior is a significant challenge. While activation steering methods, which add steering vectors to a model’s hidden states, are a promising approach, existing techniques often lack precision and interpretability in how they influence model outputs. We introduce Feature Guided Activation Additions (FGAA), a novel activation steering method that leverages insights from Contrastive Activation Addition (CAA) and Sparse Autoencoder-Targeted Steering (SAE-TS). By operating in the latent space of a Sparse Autoencoder (SAE) and employing optimization techniques to select desired SAE features, FGAA constructs precise, human-interpretable steering vectors that provide better steering effects while maintaining coherence of steered model outputs. In this regard, evaluations on Gemma-2-2B and Gemma-2-9B models across various steering tasks demonstrate that FGAA outperforms existing steering methods of CAA, SAE decoder steering, and SAE-TS. Our results also highlight important trade-offs between steering scale and general model capabilities that are consistent across all tested steering methods.
@@ 1 Introduction
The reliable and effective control of Large Language Models (LLMs) has emerged as an increasingly significant challenge in recent years. While researchers have developed various approaches to influence LLM behavior, the limitations of existing methods warrant careful consideration. Fine-tuning (Ouyang et al., 2022) offers some behavioral control but demands substantial computational resources and carefully curated datasets, making it impractical for many applications. Similarly, instruction-based approaches through prompting (Wallace et al., 2024) provide a degree of influence over model outputs but often lack robustness when faced with adversarial inputs or complex tasks. Activation steering has recently gained attention as an alternative methodology that potentially addresses these shortcomings by directly manipulating the model’s hidden state representations during inference. This technique involves introducing steering vectors at specific points in the forward pass to guide the model’s behavior in desired directions. Nevertheless, current implementations of activation steering face challenges related to interpretability, precision, and consistency which frequently resulting in unpredictable behavioral shifts and degraded output quality that limit their practical utility.
Recent work on SAE-Targeted Steering (SAE-TS) (Chalnev et al., 2024) demonstrated the value of using Sparse Autoencoders (SAEs) to extract targetable features during steering. Building on this and Contrastive Activation Addition (CAA) (Rimsky et al., 2024), we present Feature Guided Activation Additions (FGAA).
We evaluate FGAA against multiple baselines, including traditional activation steering, SAE decoder steering, and SAE-TS, across various steering tasks on both Gemma-2-2B and Gemma-2-9B models (Rivière et al., 2024). Our experiments demonstrate that FGAA achieves superior performance in both steering effectiveness and output coherence, particularly in complex steering tasks where maintaining text coherence has traditionally been challenging.
@@
    We investigate the impact of varying steering scales on the generalization capabilities of models across a diverse range of activation steering methods.
@@ 2 Related Work
Bereska and Gavves (Bereska & Gavves, 2024) outlined the central hypothesis of mechanistic interpretability: models learn human-comprehensible algorithms and can be understood, despite having no incentive to make these algorithms legible to humans during loss minimization. A key challenge in this field was identified by Scherlis et al. (Scherlis et al., 2022), who found that individual neurons often encode multiple distinct features (polysemanticity), making direct analysis of neuron behavior difficult. This is caused by superposition, the phenomenon of models representing more features than they have dimensions (Elhage et al., 2022). Sparse Autoencoders (SAEs) emerged as a solution to this challenge, with Cunningham et al. (Huben et al., 2024) demonstrating that SAEs could extract interpretable features from these superposed representations in transformer models. Bricken et al. (Bricken et al., 2023) further showed how these extracted features could be manipulated during inference to affect model behavior. Our work uses SAEs to extract interpretable features from different inputs, to construct a set of desired SAE features to steer for.
@@
Park et al. (Park et al., 2023) introduced the Linear Representation Hypothesis, showing that neural networks encode high-level concepts linearly in their representation spaces. Several studies support this hypothesis: the extraction of linear features using SAEs (Bricken et al., 2023), the effectiveness of linear probes in detecting features in the residual stream (Chanin et al., 2024), and the results from activation steering methods. We leverage this linearity assumption in both our feature selection process and its use of linear effect approximators to optimize steering vectors.
@@
Turner et al. (Turner et al., 2024) introduced activation steering (or activation engineering) to influence LLM behavior by modifying model activations during inference. Building on this work, Panickssery (Rimsky et al., 2024) introduced CAA, which computes steering vectors by averaging the difference in residual stream activations between sets of positive and negative examples of a particular behavior. Chalnev et al. (Chalnev et al., 2024) developed linear effect approximators, a linear function that predicts how steering vectors affect SAE features, allowing for targeted steering vector construction with reduced side effects. In our work, we apply the effect approximator framework to optimize CAA-derived steering vectors which are represented as SAE features.
@@ 3.2 Feature Filtering
The three filtering steps in FGAA were developed through empirical observation of feature activation patterns across multiple steering tasks. Density filtering addresses a common issue where high-density features (those that activate frequently across many inputs) tend to dominate the difference vector despite their limited task specificity. By filtering out features with activation density above $\theta=0.01$, we ensure the steering vector focuses on more specialized features that better characterize the target behavior. Similarly, BOS feature removal was implemented after observing a family of features that exclusively had the strongest activation on the BOS token (Appendix G), which often introduced artifacts in generation while contributing little to the desired steering effect. These features typically encode general linguistic patterns rather than task-specific behaviors. Finally, the selection of top $n_{1}$ positive and $n_{2}$ negative features helps eliminate noise from weakly activated features, focusing the steering vector on the most significant behavioral indicators.
@@ 3.3 Linear Approximator Optimization
We employ effect approximators (Chalnev et al., 2024) to solve for the optimal steering vector to produce the desired feature effects in $\mathbf{v}_{\text{target}}$. The linear effect approximator can be represnted as a function $\bm{\hat{y}}=\bm{x}\bm{M}+\bm{b}$, where $\bm{x}$ is the $d_{\mathrm{model}}$-dimensional steering vector, $\bm{M}$ is a $d_{\mathrm{model}}\times d_{\mathrm{sae}}$ matrix, $\bm{b}$ has dimension $d_{\mathrm{sae}}$, and $\bm{\hat{y}}$ is the predicted steering effects vector of dimension $d_{\mathrm{sae}}$.
@@ 4.1 Effectiveness of FGAA for Steering
For our evaluations, FGAA is implemented using a pre-trained Gemma Scope (Lieberum et al., 2024) SAE with 16,384 features for the residual stream at layer 12 for Gemma-2-2B and Gemma-2-9B models. We selected these two models due to both computational constraints and the availability of open pre-trained SAE weights. Similarly, we apply steering to the residual stream at layer 12 and utilize pretrained effect approximators from (Chalnev et al., 2024) for both Gemma models. We focus on layer 12 in our evaluation, as collecting training data for effect approximators is time-intensive and must be done separately for each layer. Additionally, only layer 12 approximators for the models above have been made publicly available.
We evaluate FGAA against existing steering methods using the evaluation framework from (Chalnev et al., 2024), employing gpt-4o-mini to assess both behavioral alignment and coherence on a 1-10 scale, which we then rescale to the range [0,1]. Let $B$ represent the behavioral score which measures steering target achievement, and $C$ represent coherence which evaluates semantic correctness post-steering (exact criterion in Appendix C). We define the Behavioral-Coherence Score (BCS) as:
@@
Goal CAA SAE SAE-TS FGAA (Ours) CAA SAE SAE-TS FGAA (Ours)
@@
Figure 3: Plots showing mean BCS with 95% confidence intervals for the CAA, SAE, SAE-TS and FGAA steering methods on 9 tasks, for Gemma-2-9B.
@@
  • Programmatic Feature Selection: SAE-TS and SAE methods requires manual selection of a single feature to steer towards. FGAA programmatically identifies a spectrum of relevant features, while preserving the relationships in magnitude between them (refer to Table A.1 for an example). This is more realistic as, especially in lower width SAEs, it cannot be expected that every concept the LLM learns be cleanly encoded as an SAE latent. The presence of polysemantic and uninterpretable features extracted from SAEs across varying widths and models shows strong evidence for this, prompting research into Meta-SAEs (Anonymous, 2025) to further break down superposition. Instead, by representing concepts as a target vector in the feature space, we are able to achieve more precise concept representation. In larger width SAEs, this automated feature selection becomes more helpful due to the phenomena of feature splitting (Chanin et al., 2024), where a feature represented in a single latent in a smaller SAE can split into two or more latents in a larger SAE. FGAA systematically handles such cases by programmatically determine the relative steering magnitudes between semantically similar features. FGAA also handles the rare case where only targeting a single feature is the most effective steering approach, as detailed in Appendix D.
  • Interpretability: While current CAA methods operate in opaque activation spaces, FGAA’s backwards approach—determining desired effects in feature space before constructing steering vectors—provides explicit control over which features are steered, and to what extent. Through automatic interpertability (Paulo et al., 2024), SAE features can be labelled with human-interpretable descriptions (examples in Appendix B), allowing practitioners to directly understand which semantic aspects of the model’s behavior are being modified during steering. This transparency also allows us to filter away redundant components of the steering vector (via methods in Section 3.2) which would otherwise be present in CAA-derived vectors, allowing for more precise steering interventions.
@@ 4.2 Effects of Steering on General Model Capabilities
We evaluate the impact of steering methods on model capabilities through perplexity testing on the OpenWebText (Gokaslan & Cohen, 2019) dataset and performance on MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024) benchmarks. MMLU is a comprehensive evaluation benchmark that tests AI models using multiple choice questions spanning 57 different subjects, from STEM fields to humanities and social sciences. While the original MMLU primarily focuses on testing factual knowledge, MMLU-Pro builds upon this foundation by introducing more complex questions that require deeper reasoning abilities and increases the number of possible answers from 4 to 10 per question.
@@
This degradation pattern is further corroborated by benchmark performance on MMLU and MMLU-Pro (Figures 5(a), 5(b)). Both benchmarks demonstrate that model capabilities are largely preserved at lower steering scales but deteriorate as steering intensity increases. At scales below 50, all methods maintain close to baseline performance. However, beyond this threshold, we observe a consistent pattern of degradation across all steering approaches, with performance declining sharply between scales of 50 and 150 before converging near zero at higher scales.
@@
An intriguing observation is the slight increase in MMLU-Pro performance at low steering scales for CAA, SAE-TS, and SAE methods. This phenomenon may be analogous to how low levels of noise can enhance LLM inference performance, similar to effects observed with techniques like NEFTune (Jain et al., 2024). At very low steering scales, these steering vectors might function as beneficial noise that temporarily improves model capabilities before the more disruptive effects of steering become dominant at higher scales. The absence of this initial performance bump in FGAA, which instead shows stable performance, suggests its steering interventions are more precisely targeted. This aligns with FGAA’s design objective of creating focused steering interventions through feature space optimization rather than introducing broader activation perturbations. While this observation merits further investigation to fully understand the underlying mechanisms, such analysis falls outside the scope of this paper.
@@ 6 Future Work
We believe the most promising direction to pursue would be applying FGAA to existing works in the activation steering space, to see if FGAA performance improvements carry over to safety tasks such as controlling sycophancy, hallucination and refusal in RLHF models (Rimsky et al., 2024) and reducing their social biases (Durmus et al., 2024).
@@ 7 Conclusion
Our analysis revealed important insights about activation steering in general: performance degrades notably above certain steering scales, and there exists a fundamental tradeoff between steering strength and preservation of model capabilities.
@@ 8 Acknowledgements
We would like to express our sincere gratitude to our research mentors, Dr Tan Guoxian and Dr Yan Ming, for their guidance throughout our research. We are grateful to Mr Chan Kwang Wen for contributing OpenAI API credits that enabled our evaluations. Special thanks to Chen Guang, Mr Slava Chalnev and Mr Logan Riggs for their insightful discussions on SAEs and activation steering. We also thank Chen Guang for providing the necessary compute resources for our work.
@@ Appendix B Examples of Constructed Filtered Target Vectors
"The media is spreading disinformation to manipulate
    public opinion"
"Whistleblowers are being silenced to cover up a
    massive scandal"
@@
"Peer-reviewed studies provide reliable,
evidence-based conclusions"
@@
"<bos>I think this is a photo of a giant squid
    attacking a Russian submarine,
    and it is one of the most Incredible Aliens captured
    in Antarctica! These mind"
@@
    Program called the Advanced Space Agency (AASPA)
    and the US"
@@
    our Universe. But, there may be Alien Lifeforms
    on Mars that could be Behind The Pyramids on Mars. "
@@
    of Extraterrestrial Technology to the World.
    Recently, a group of British researchers
    have claimed that they have uncovered"
"<bos>I think the FBI has captured
    a ‘Super-Being’ named Ra from an Alien Craft
    that fell to Earth 6,000 years Ago, according"
"<bos>I think that we are being manipulated by
    Extraterrestrial Races through The Vatican
    and the Catholic Church, so I began an ambitious
    Project to uncover the truth. "
@@
"I’m devoted to you and cherish every moment we spend
    together"
@@
"I’m passionate about our relationship and excited for our
    future"
@@
"<bos>I think every girl and woman knows what it feels
    like to have a crush. At one point or another, we
    have all been loved so hard by a man"
@@
"<bos>I think this has to be the most amazing
    and inspirational story. I love my parents,
    and there is no one more special to me. The best"
"<bos>I think it’s so important to capture these
    little moments with my love. I think she is
    so beautiful. She is 1 week older than my sweet"
"<bos>I think it’s safe to say that we simply love
    all of Justin and Hailey’s wedding photos
    and it’s no wonder why we always swoon"
"<bos>I think the best part about dating is being your
    own best friend. I wish I knew how much I loved you
    in a different way, and the way"